# Understanding and Benchmarking the Trustworthiness in Multimodal LLMs for Video Understanding

### v2

@@ Abstract
Recent advancements in multimodal large language models for video understanding (videoLLMs) have enhanced their capacity to process complex spatiotemporal data. However, challenges such as factual inaccuracies, harmful content, biases, hallucinations, and privacy risks compromise their reliability. This study introduces Trust-videoLLMs, a first comprehensive benchmark evaluating 23 state-of-the-art videoLLMs (5 commercial, 18 open-source) across five critical dimensions: truthfulness, robustness, safety, fairness, and privacy. Comprising 30 tasks with adapted, synthetic, and annotated videos, the framework assesses spatiotemporal risks, temporal consistency and cross-modal impact. Results reveal significant limitations in dynamic scene comprehension, cross-modal perturbation resilience and real-world risk mitigation. While open-source models occasionally outperform, proprietary models generally exhibit superior credibility, though scaling does not consistently improve performance. These findings underscore the need for enhanced training datat diversity and robust multimodal alignment. Trust-videoLLMs provides a publicly available, extensible toolkit for standardized trustworthiness assessments, addressing the critical gap between accuracy-focused benchmarks and demands for robustness, safety, fairness, and privacy. Code are available at: https://github.com/wangyouze/Trust-videoLLMs
@@ 1 Introduction
Recent advancements in video large language models demonstrate their superior capability to process dynamic visual information across diverse multimodal benchmarks [1, 2, 3, 4, 5, 6], positioning them as foundational models for understanding the temporal and spatial complexities of real-world multimodal data. These developments mark significant progress toward artificial general intelligence. However, despite efforts to align with human values, videoLLMs face critical trustworthiness challenges, including factual inaccuracies [1, 7], temporal inconsistency [8], harmful content generation [9], biases [10], hallucinations [4, 11], and privacy vulnerabilities [12]. The inherent spatiotemporal complexities of video data intensify these problems, compromising the dependability of videoLLMs and generating widespread apprehension across academic circles, governance bodies, and civil society.
@@
Video-MME [1] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 6 13(4)
TemporalBench [13] $\times$ $\times$ $\times$ $\times$ $\times$ 6 15(4)
VELOCITI [14] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 7 8(3)
TempCompass [8] $\times$ $\times$ $\times$ $\times$ $\times$ 4 11(1)
Video-Bench [7] $\times$ $\times$ $\times$ $\times$ $\times$ 10 8(0)
MVBench [7] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 20 14(1)
VideoHallucer [4] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 5 9(2)
HAVEN [11] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 4 12(2)
Video-SafetyBench [6] $\times$ $\times$ $\times$ $\times$ $\times$ 13 24(7)
SafeVidBench [15] $\times$ $\times$ $\times$ $\times$ $\times$ 7 16(6)
@@
Compared to image-based MLLMs [16, 17, 18, 19], which process static visual data with limited exposure to external interference, videoLLMs excel in multimodal comprehension by integrating complex temporal and spatial interactions between visual, auditory, and textual inputs. Current multimodal large language models (MLLMs) benchmarks, covering both image- and video-based evaluations, primarily assess video understanding accuracy [7, 1] and long-video comprehension reliability [20] but often overlook critical dimensions such as adversarial robustness, safety, fairness, and privacy. Concurrent work [6, 15] focus on the safety dimension of videoLLMs. Image-based MLLMs benchmark [21], designed for static visual tasks, are inadequate to address trustworthiness risks arising from the dynamic nature of video content, necessitating comprehensive benchmarks tailored to the spatiotemporal complexities of videoLLMs (The limitations of these benchmarks are detailed in Table I).
This study presents Trust-videoLLMs, a comprehensive framework for evaluating the trustworthiness of MLLMs for video understanding and analysis, as shown in Figure 1. Extending the five core dimensions of the TrustLLM [22]: truthfulness, safety, robustness, fairness, and privacy, we introduce novel tasks tailored to the spatiotemporal dynamics and multimodal nature of video data. A multi-level evaluation approach examines the evolution of multimodal risks in dynamic scenarios and the cross-modal impact of temporal visual inputs on foundational language models, revealing critical vulnerabilities in videoLLMs. The evaluation system comprises 30 tasks, including: (1) spatiotemporal tasks different from image-based trustworthiness benchmarks to establish dynamic scenario standards; (2) analysis of multimodal input interactions affecting videoLLMs decisions; and (3) assessment of model robustness in core video tasks and safety risks in real-world applications. To support this, we construct a large-scale dataset integrating task-adapted existing datasets, synthetic data generated using advanced text/image-to-video tools (e.g., Kling  [^1] https://app.klingai.com/cn/, Jimeng  [^2] https://jimeng.jianying.com/), and manually collect and annotated data, ensuring diverse scenario coverage.
@@
  • While open-source videoLLMs occasionally outperform proprietary models on specific truthfulness benchmarks [1, 3, 4], their overall trustworthiness remains lower than that of mainstream proprietary models. Our evaluation reveals that Claude series and Gemini1.5 series demonstrate superior security and trustworthiness.
  • Larger parameter counts don’t consistently translate to better performance. Instead, architectural design and training strategies play a more critical role. Notably, the 7B-paramete Qwen2.5-VL achieves strong performance across multiple dimensions, ranking 9th overall and outperforming many larger models.
@@
T.2 Temporal Reasoning QA [8] Accuracy Dis.
T.3 Video Description [23, 8] LLLM-score, score, B, M, C, R Gen.
T.4 Events Understanding [24] Accuracy Dis.
T.5 Hallucination in Videos [4] Accuracy Dis.
R.1 OOD Videos Captioning [25] LLLM-score, B, M, C, R Gen.
R.2 Noise Videos QA [3] Acc Dis.
R.3 Temporal Understanding Consistency [3] Accuracy Dis.
R.4 Adversarial Attack for Classification [3] Accuracy Dis.
R.5 Adversarial Attack for Captioning [3] LLLM-score, B, M, C, R Gen.
@@
S.2 NSFW Prompts Execution [27] RtA,LLLM-score,Toxicity-score Gen.
@@
S.6 Deepfake Identification [30] Accuracy Dis.
S.7 Jailbreak Attacks [31, 32, 9] RtA, LLLM-score, Toxicity-score Gen.
F.1 Stereotype Impact generation [23] LLLM-score Gen.
@@
P.1 Privacy content Recognition [33] Accuracy, Precision,Recall, F1 Dis.
@@ 2.1 Multimodal Understanding Benchmarks
Significant efforts have been devoted to developing evaluation frameworks to advance MLLMs. Existing studies assess the trustworthiness of image-based MLLMs across multiple dimensions. For example, POPE [35] systematically evaluates hallucination issues, while SafeBench [36] and MM-SafetyBench [32] provide specialized benchmarks for safety alignment in multimodal scenarios. RTVLM [37] introduces a red-teaming dataset evaluating truthfulness, safety, privacy, and fairness. MultiTrust [21] offers a comprehensive toolkit assessing: truthfulness, robustness, safety, fairness, and privacy. However, these benchmarks primarily address static image data and text prompts, failing to capture the spatiotemporal complexities of videoLLMs.
@@ 2.2 Trustworthiness Evaluation on Video Understanding
Video understanding, a key branch of MLLMs, demonstrate superior capabilities in processing dynamic visual information and diverse multimodal benchmarks [1, 2, 3, 4, 5, 6], establishing them as foundational models for understanding the spatiotemporal complexities of real-world multimodal data. However, despite efforts to align videoLLMs with human values, they face critical trustworthiness challenges [7, 8, 4, 12]. Compared to image-based MLLMs (e.g., LLaVA-Next [17]), videoLLMs exhibit enhanced multimodal understanding by integrating complex spatiotemporal interactions across visual, auditory, and textual inputs. However, current evaluation benchmarks for videoLLMs primarily focus on video understanding accuracy [7, 1] and long-video comprehension reliability [20]. Concurrent studies [6, 15] mainly focus on the safety dimension of videoLLMs. Existing image-based MLLM benchmarks [21], designed for static visual tasks, are ill-equipped to address trustworthiness risks arising from the dynamic nature of video content. Consequently, there is an urgent need for a comprehensive evaluation benchmark tailored to the spatiotemporal complexities of videoLLMs from truthfulness, robustness, safety, fairness, and privacy, enabling a systematic and holistic assessment of their trustworthiness to guide model optimization and ensure reliable deployment.
@@ 3.1.1 Truthfulness in videoLLMs: Assessing Accuracy and Reliability
Truthfulness is vital for ensuring that videoLLMs provide accurate and reliable outputs based on dynamic visual content. Unlike static images, videos require reasoning over temporal sequences, increasing the risk of errors and hallucinations. Departing from prior work [7, 8, 1], we assess truthfulness through two key aspects:
@@ 3.1.2 Robustness in videoLLMs: Assessing Consistency and Resistance
OOD Robustness (O.) evaluates a videoLLM’s generalization to unseen domains, including diverse video types and natural noise. We utilize an OOD video dataset [25] and natural noise to assess model robustness under OOD conditions (Tasks R.1 and R.2).
@@
Adversarial Robustness (A.) evaluates susceptibility to adversarial inputs—a well-known vulnerability of deep neural networks [38]. We assess the model’s resilience to perturbations crafted to mislead video understanding, using keyframes selected via uniform sampling and optical flow analysis. Adversarial examples are generated with the MI-CWA algorithm [39], and performance is evaluated on video classification (Task R.4) and captioning (Task R.5).
@@ 3.1.3 Safety in videoLLMs: Assessing Output Security
Toxicity in Generated Content (G.). Toxicity refers to outputs containing pornography, violence, blood, or hate speech. We assess videoLLMs’ ability to detect and describe NSFW content using videos from the BigPorn, Violence, and HateSpeech datasets (Task S.1). Rejection rates under harmful instructions are evaluated using HarmBench [27] (Task S.2). To examine how video context influences toxic responses, prompts from RealToxicityPrompts [28] are paired with semantically related and unrelated videos across five toxicity categories (Task S.3).
@@
Safety Against Malicious Manipulations (S.). Robustness against adversarial manipulation is vital. We assess deepfake detection using manipulated videos (Task S.6) and evaluate resistance to jailbreak attacks (Task S.7). This includes two image-based methods—FigStep [31] and MMsafetyBench [32]—converted to video, and one native video-based attack, VideoJail [9].
@@ 3.1.4 Fairness and Bias in VideoLLMs: Assessing Equity and Bias
Bias from Data-Driven Influences (B.). VideoLLMs trained on large-scale datasets may inherit demographic biases, potentially generating stereotypical outputs. We evaluate bias manifestation through three complementary methods: analyzing stereotype presence using videos from OpenVid-1M [23] covering occupation, gender, age, and race with targeted prompts (Task F.1); employing established text-based bias benchmarks [40, 41] to generate videos for stereotype classification (Task F.2); and extending bias evaluation to scenarios with different attributes description including age, gender, and skin tone (Task F.3). These assessments evaluate model bias awareness arising from data-driven influences.
@@ 3.1.5 Privacy in VideoLLMs: Assessing Information Protection and Inference Control
Privacy Awareness (R.). This component evaluates whether videoLLMs can detect and appropriately respond to privacy-sensitive content, such as faces, license plates, and ID cards. We first assess identification capabilities using videos from BIV-Priv [33], which include items like passports, credit cards, and private letters (Task P.1). We further evaluate privacy perception using real-world YouTube videos containing transient private information, such as phone or computer screens, license plates, and delivery addresses by QA method (Task P.2).
Control Over Privacy Inference (I.). VideoLLMs should avoid inferring or generating private information, especially without explicit prompts. We adapt the InfoFlow Expectation task [34] to multimodal settings, pairing it with videos to assess model agreement on privacy use (Task P.3). To evaluate protection of celebrity privacy, we use videos from diverse domains (sports, entertainment, politics, music), examining whether the model refrains from disclosing personal details (Task P.4). We also test the tendency to infer implicit privacy-sensitive content using videos from the OpenVid-1M dataset [23] (Task P.5).

### v3

# Benchmarking Trustworthiness in Multimodal LLMs for Video Understanding

@@ Abstract
Recent advancements in multimodal large language models for video understanding (videoLLMs) have enhanced their capacity to process complex spatiotemporal data. However, challenges such as factual inaccuracies, harmful content, biases, hallucinations, and privacy risks compromise their reliability. This study introduces Trust-videoLLMs, a first comprehensive benchmark evaluating 23 state-of-the-art videoLLMs (5 commercial, 18 open-source) across five critical dimensions: truthfulness, robustness, safety, fairness, and privacy. Comprising 30 tasks with adapted, synthetic, and annotated videos, the framework assesses spatiotemporal risks, temporal consistency and cross-modal impact. Results reveal significant limitations in dynamic scene comprehension, cross-modal perturbation resilience and real-world risk mitigation. While open-source models occasionally outperform, proprietary models generally exhibit superior credibility, though scaling does not consistently improve performance. These findings underscore the need for enhanced training data diversity and robust multimodal alignment. Trust-videoLLMs provides a publicly available, extensible toolkit for standardized trustworthiness assessments, addressing the critical gap between accuracy-focused benchmarks and demands for robustness, safety, fairness, and privacy. Code are available at: https://github.com/wangyouze/Trust-videoLLMs
[…]
@@ 1 Introduction
Recent advancements in video large language models demonstrate their superior capability to process dynamic visual information across diverse multimodal benchmarks [fu2024video, li2025temporal, li2024mvbench, wang2024videohallucer, guan2024hallusionbench, liu2025video], positioning them as foundational models for understanding the temporal and spatial complexities of real-world multimodal data. These developments mark significant progress toward artificial general intelligence. However, despite efforts to align with human values, videoLLMs face critical trustworthiness challenges, including factual inaccuracies [fu2024video, ning2023video], temporal inconsistency [liu2024tempcompass], harmful content generation [hu2025videojail], biases [park2025assessing], hallucinations [wang2024videohallucer, gao2025exploring], and privacy vulnerabilities [tang2025video]. The inherent spatiotemporal complexities of video data intensify these problems, compromising the dependability of videoLLMs and generating widespread apprehension across academic circles, governance bodies, and civil society.
@@
Video-MME [fu2024video] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 6 13(4)
TemporalBench [cai2024temporalbench] $\times$ $\times$ $\times$ $\times$ $\times$ 6 15(4)
VELOCITI [saravanan2024velociti] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 7 8(3)
TempCompass [liu2024tempcompass] $\times$ $\times$ $\times$ $\times$ $\times$ 4 11(1)
Video-Bench [ning2023video] $\times$ $\times$ $\times$ $\times$ $\times$ 10 8(0)
MVBench [ning2023video] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 20 14(1)
VideoHallucer [wang2024videohallucer] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 5 9(2)
HAVEN [gao2025exploring] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ 4 12(2)
Video-SafetyBench [liu2025video] $\times$ $\times$ $\times$ $\times$ $\times$ 13 24(7)
SafeVidBench [wang2025safevid] $\times$ $\times$ $\times$ $\times$ $\times$ 7 16(6)
@@
Compared to image-based MLLMs [zhu2023minigpt, liu2024llavanext, Qwen2.5-VL, hurst2024gpt], which process static visual data with limited exposure to external interference, videoLLMs excel in multimodal comprehension by integrating complex temporal and spatial interactions between visual, auditory, and textual inputs. Current multimodal large language models (MLLMs) benchmarks, covering both image- and video-based evaluations, primarily assess video understanding accuracy [ning2023video, fu2024video] and long-video comprehension reliability [wang2024lvbench] but often overlook critical dimensions such as adversarial robustness, safety, fairness, and privacy. Concurrent work [liu2025video, wang2025safevid] focus on the safety dimension of videoLLMs. Image-based MLLMs benchmark [zhang2024benchmarking], designed for static visual tasks, are inadequate to address trustworthiness risks arising from the dynamic nature of video content, necessitating comprehensive benchmarks tailored to the spatiotemporal complexities of videoLLMs (The limitations of these benchmarks are detailed in Table I).
This study presents Trust-videoLLMs, a comprehensive framework for evaluating the trustworthiness of MLLMs for video understanding and analysis, as shown in Figure 1. Extending the five core dimensions of the TrustLLM [sun2024trustllm]: truthfulness, safety, robustness, fairness, and privacy, we introduce novel tasks tailored to the spatiotemporal dynamics and multimodal nature of video data. A multi-level evaluation approach examines the evolution of multimodal risks in dynamic scenarios and the cross-modal impact of temporal visual inputs on foundational language models, revealing critical vulnerabilities in videoLLMs. The evaluation system comprises 30 tasks, including: (1) spatiotemporal tasks different from image-based trustworthiness benchmarks to establish dynamic scenario standards; (2) analysis of multimodal input interactions affecting videoLLMs decisions; and (3) assessment of model robustness in core video tasks and safety risks in real-world applications. To support this, we construct a large-scale dataset integrating task-adapted existing datasets, synthetic data generated using advanced text/image-to-video tools (e.g., Kling  [^1] https://app.klingai.com/cn/, Jimeng  [^2] https://jimeng.jianying.com/), and manually collect and annotated data, ensuring diverse scenario coverage.
@@
  • While open-source videoLLMs occasionally outperform proprietary models on specific truthfulness benchmarks [fu2024video, li2024mvbench, wang2024videohallucer], their overall trustworthiness remains lower than that of mainstream proprietary models. Our evaluation reveals that Claude series and Gemini1.5 series demonstrate superior security and trustworthiness.
  • Larger parameter counts don’t consistently translate to better performance. Instead, architectural design and training strategies play a more critical role. Notably, the 7B-parameter Qwen2.5-VL achieves strong performance across multiple dimensions, ranking 9th overall and outperforming many larger models.
@@
T.2 Temporal Reasoning QA [liu2024tempcompass] Accuracy Dis.
T.3 Video Description [nan2024openvid, liu2024tempcompass] LLLM-score, score, B, M, C, R Gen.
T.4 Events Understanding [ZhXuCoAAAI18] Accuracy Dis.
T.5 Hallucination in Videos [wang2024videohallucer] Accuracy Dis.
R.1 OOD Videos Captioning [bashmal2023capera] LLLM-score, B, M, C, R Gen.
R.2 Noise Videos QA [li2024mvbench] Acc Dis.
R.3 Temporal Understanding Consistency [li2024mvbench] Accuracy Dis.
R.4 Adversarial Attack for Classification [li2024mvbench] Accuracy Dis.
R.5 Adversarial Attack for Captioning [li2024mvbench] LLLM-score, B, M, C, R Gen.
@@
S.2 NSFW Prompts Execution [mazeika2024harmbench] RtA,LLLM-score,Toxicity-score Gen.
@@
S.6 Deepfake Identification [dufour2019deepfakes] Accuracy Dis.
S.7 Jailbreak Attacks [gong2025figstep, liu2024mm, hu2025videojail] RtA, LLLM-score, Toxicity-score Gen.
F.1 Stereotype Impact generation [nan2024openvid] LLLM-score Gen.
@@
P.1 Privacy content Recognition [sharma2023disability] Accuracy, Precision,Recall, F1 Dis.
@@ 2.1 Multimodal Understanding Benchmarks
Significant efforts have been devoted to developing evaluation frameworks to advance MLLMs. Existing studies assess the trustworthiness of image-based MLLMs across multiple dimensions. For example, POPE [li2023evaluating] systematically evaluates hallucination issues, while SafeBench [ying2024safebench] and MM-SafetyBench [liu2024mm] provide specialized benchmarks for safety alignment in multimodal scenarios. RTVLM [li2024red] introduces a red-teaming dataset evaluating truthfulness, safety, privacy, and fairness. MultiTrust [zhang2024benchmarking] offers a comprehensive toolkit assessing: truthfulness, robustness, safety, fairness, and privacy. However, these benchmarks primarily address static image data and text prompts, failing to capture the spatiotemporal complexities of videoLLMs.
@@ 2.2 Trustworthiness Evaluation on Video Understanding
Video understanding, a key branch of MLLMs, demonstrate superior capabilities in processing dynamic visual information and diverse multimodal benchmarks [fu2024video, li2025temporal, li2024mvbench, wang2024videohallucer, guan2024hallusionbench, liu2025video], establishing them as foundational models for understanding the spatiotemporal complexities of real-world multimodal data. However, despite efforts to align videoLLMs with human values, they face critical trustworthiness challenges [ning2023video, liu2024tempcompass, wang2024videohallucer, tang2025video]. Compared to image-based MLLMs (e.g., LLaVA-Next [liu2024llavanext]), videoLLMs exhibit enhanced multimodal understanding by integrating complex spatiotemporal interactions across visual, auditory, and textual inputs. However, current evaluation benchmarks for videoLLMs primarily focus on video understanding accuracy [ning2023video, fu2024video] and long-video comprehension reliability [wang2024lvbench]. Concurrent studies [liu2025video, wang2025safevid] mainly focus on the safety dimension of videoLLMs. Existing image-based MLLM benchmarks [zhang2024benchmarking], designed for static visual tasks, are ill-equipped to address trustworthiness risks arising from the dynamic nature of video content. Consequently, there is an urgent need for a comprehensive evaluation benchmark tailored to the spatiotemporal complexities of videoLLMs from truthfulness, robustness, safety, fairness, and privacy, enabling a systematic and holistic assessment of their trustworthiness to guide model optimization and ensure reliable deployment.
@@ 3.1.1 Truthfulness in videoLLMs: Assessing Accuracy and Reliability
Truthfulness is vital for ensuring that videoLLMs provide accurate and reliable outputs based on dynamic visual content. Unlike static images, videos require reasoning over temporal sequences, increasing the risk of errors and hallucinations. Departing from prior work [ning2023video, liu2024tempcompass, fu2024video], we assess truthfulness through two key aspects:
@@ 3.1.2 Robustness in videoLLMs: Assessing Consistency and Resistance
OOD Robustness (O.) evaluates a videoLLM’s generalization to unseen domains, including diverse video types and natural noise. We utilize an OOD video dataset [bashmal2023capera] and natural noise to assess model robustness under OOD conditions (Tasks R.1 and R.2).
@@
Adversarial Robustness (A.) evaluates susceptibility to adversarial inputs—a well-known vulnerability of deep neural networks [szegedy2013intriguing]. We assess the model’s resilience to perturbations crafted to mislead video understanding, using keyframes selected via uniform sampling and optical flow analysis. Adversarial examples are generated with the MI-CWA algorithm [chen2023rethinking], and performance is evaluated on video classification (Task R.4) and captioning (Task R.5).
@@ 3.1.3 Safety in videoLLMs: Assessing Output Security
Toxicity in Generated Content (G.). Toxicity refers to outputs containing pornography, violence, blood, or hate speech. We assess videoLLMs’ ability to detect and describe NSFW content using videos from the BigPorn, Violence, and HateSpeech datasets (Task S.1). Rejection rates under harmful instructions are evaluated using HarmBench [mazeika2024harmbench] (Task S.2). To examine how video context influences toxic responses, prompts from RealToxicityPrompts [gehman2020realtoxicityprompts] are paired with semantically related and unrelated videos across five toxicity categories (Task S.3).
@@
Safety Against Malicious Manipulations (S.). Robustness against adversarial manipulation is vital. We assess deepfake detection using manipulated videos (Task S.6) and evaluate resistance to jailbreak attacks (Task S.7). This includes two image-based methods—FigStep [gong2025figstep] and MMsafetyBench [liu2024mm]—converted to video, and one native video-based attack, VideoJail [hu2025videojail].
@@ 3.1.4 Fairness and Bias in VideoLLMs: Assessing Equity and Bias
Bias from Data-Driven Influences (B.). VideoLLMs trained on large-scale datasets may inherit demographic biases, potentially generating stereotypical outputs. We evaluate bias manifestation through three complementary methods: analyzing stereotype presence using videos from OpenVid-1M [nan2024openvid] covering occupation, gender, age, and race with targeted prompts (Task F.1); employing established text-based bias benchmarks [nadeem2020stereoset, nangia2020crows] to generate videos for stereotype classification (Task F.2); and extending bias evaluation to scenarios with different attributes description including age, gender, and skin tone (Task F.3). These assessments evaluate model bias awareness arising from data-driven influences.
@@ 3.1.5 Privacy in VideoLLMs: Assessing Information Protection and Inference Control
Privacy Awareness (R.). This component evaluates whether videoLLMs can detect and appropriately respond to privacy-sensitive content, such as faces, license plates, and ID cards. We first assess identification capabilities using videos from BIV-Priv [sharma2023disability], which include items like passports, credit cards, and private letters (Task P.1). We further evaluate privacy perception using real-world YouTube videos containing transient private information, such as phone or computer screens, license plates, and delivery addresses by QA method (Task P.2).
Control Over Privacy Inference (I.). VideoLLMs should avoid inferring or generating private information, especially without explicit prompts. We adapt the InfoFlow Expectation task [mireshghallah2023can] to multimodal settings, pairing it with videos to assess model agreement on privacy use (Task P.3). To evaluate protection of celebrity privacy, we use videos from diverse domains (sports, entertainment, politics, music), examining whether the model refrains from disclosing personal details (Task P.4). We also test the tendency to infer implicit privacy-sensitive content using videos from the OpenVid-1M dataset [nan2024openvid] (Task P.5).